11) A regression of income (in thousands of dollars) on education (in years) is performed on data collected from 27 individuals. It had an intercept of 31.8 and a slope of 6.1. Answer the following questions:

a) Fill in the missing information in the ANOVA table.

Sum of Squares

df

Mean Squares

F

Regression

2,000

Residual

Total

20,000

b) What is the predicted income for someone with 11 years of education?

c) what is the coefficient of determination?

d) Using the context of the problem, explain the meaning of the slope being 6.1.
Answer #1

Ss (Residual) SS(Residual n=27 = 20,000 - 2000 1800o If (Regression) = -2=27-2=25 afl Residual = 2. M.S. (R) ss 80 2000 25 df